### v3

@@ 3 Methodology
3.1 Problem statement and two solutions
The objective of building damage assessment is to interpret EO data covering the affected area and generate a building damage map that reflects the extent of damage to buildings. To achieve this, two common solutions are typically employed.
One is to directly treat the building damage assessment task as a single semantic segmentation task (Adriano et al., 2021; Gupta and Shah, 2021). The pre- and post-event images are taken as inputs of the model, and then the final damage map is directly predicted. This process can be formalized as $\mathbf{Y}^{dam}=\mathcal{M}^{seg}(\mathbf{X}^{\mathrm{T}_{1}},\mathbf{X}^{\mathrm{T}_{2}})$, where is $\mathbf{X}^{\mathrm{T}_{1}}$ is the pre-event imagery, $\mathbf{X}^{\mathrm{T}_{2}}$ is the post-event imagery, $\mathcal{M}^{seg}\left(\cdot\right)$ is a semantic segmentation model, $\mathbf{Y}^{dam}$ is the obtained damage map. In the context of this paper, $\mathbf{X}^{\mathrm{T}_{1}}$ is VHR optical imagery and $\mathbf{X}^{\mathrm{T}_{2}}$ is VHR SAR imagery.
The second solution adopts the task decoupling approach (Gupta et al., 2019; Zheng et al., 2021), which breaks down building damage assessment into two subtasks: the building localization task, i.e., separating the building from the background, and the damage classification task, i.e., focusing on the classification between different levels of damage. This solution can be formulated as $\mathbf{Y}^{loc}=\mathcal{M}^{loc}(\mathbf{X}^{\mathrm{T}_{1}})$ and $\mathbf{Y}^{clf}=\mathcal{M}^{clf}(\mathbf{X}^{\mathrm{T}_{1}},\mathbf{X}^{\mathrm{T}_{2}})$, where $\mathbf{Y}^{loc}$ is the building localization map, $\mathbf{Y}^{clf}$ is the damage classification map, $\mathcal{M}^{loc}\left(\cdot\right)$ and $\mathcal{M}^{clf}\left(\cdot\right)$ are models for building localization and damage classification tasks, respectively. $\mathcal{M}^{loc}\left(\cdot\right)$ and $\mathcal{M}^{clf}\left(\cdot\right)$ can be two separate models (Gupta et al., 2019) or a unified multi-task learning model (Zheng et al., 2021; Chen et al., 2022a, 2024). The final building damage map is obtained by combining the two outputs using a simple mask operation: $\mathbf{Y}^{dam}=\mathbf{Y}^{loc}\odot\mathbf{Y}^{clf}$.
Since the aim of this article is not only to provide a large-scale multimodal dataset to support all-weather disaster response but also to offer insights for designing appropriate methods in future research, we will employ and compare both solutions in our experiments.
@@ 3.2 Benchmark suites
We evaluate several advanced deep network architectures from both the computer vision and EO communities on Bright. Since the building damage assessment task can be considered a specialized semantic segmentation task, we selected two well-known segmentation networks from the computer vision field: UNet (Ronneberger et al., 2015) and DeepLabV3+ (Chen et al., 2018). Additionally, we included several state-of-the-art networks from the EO community: SiamAttnUNet (Adriano et al., 2021), SiamCRNN (Chen et al., 2020), ChangeOS (Zheng et al., 2021), DamageFormer (Chen et al., 2022a), and ChangeMamba (Chen et al., 2024). These networks encompass a broad range of representative DL architectures, including convolutional neural networks (CNNs), recurrent neural networks (RNNs), Transformers, and the more recent Mamba architecture.
Among these seven networks, UNet, DeepLabV3+, and SiamAttnUNet adopt the first solution, directly treating building damage assessment as a single semantic segmentation task. In contrast, SiamCRNN, ChangeOS, DamageFormer, and ChangeMamba adopt the second solution by decoupling the task into building localization and damage classification tasks.
Beyond supervised deep learning, Bright also enables the evaluation of broader strategies commonly explored in EO and vision communities:
@@
Table 5: Accuracy assessment for different DL models on the test set under the standard machine learning division (set-level mIoU). The highest values are highlighted in purple, and the second-highest results are highlighted in teal.
[…]
@@ 3.3 Model training
To train these supervised models, we use a combination of cross-entropy loss and Lovasz softmax loss (Berman et al., 2018). Cross-entropy loss serves as the basic loss function for dense prediction tasks, while Lovasz softmax loss effectively addresses sample imbalance between non-building and building pixels and across different damage levels.
For UNet, DeepLabV3+, and SiamAttnUNet, which directly predict damage maps from the input multimodal image pairs, the training loss function is defined as:
@@
For zero-shot and one-shot setups, we incorporate UDA and SSL techniques to better utilize unlabeled and limited target disaster data. The specifics of these training adjustments, along with the UMCD and UMIM approaches, are provided in Appendix E, F, G, and H, respectively.
@@
3.4 Accuracy assessment
We adopt overall accuracy (OA), F1 score (F1), and mean intersection over union (mIoU) to evaluate the performance of DL models on building damage assessment. These metrics are commonly used in previous research on building damage assessment (Zheng et al., 2021). Following the setup in prior unimodal building damage assessment studies and the related xView2 Challenge (Gupta et al., 2019), the F1 score is used to assess the performance of DL models in the building localization and damage classification subtasks. Finally, OA and mIoU are employed to measure the overall quality of the building damage map, providing a comprehensive assessment of the models’ ability to localize buildings and classify damage levels accurately.
4 Results and Analysis
4.1 Evaluation results on standard machine learning split
@@
Figure 6: Damage maps predicted by different models on the test images of 14 events. The meaning of the color in reference maps and damage maps is consistent with Figure 4.
Finally, Figure 6 shows some of the building damage maps obtained by the seven models on the test set.
4.2 What have models learned and what can they learn?
Figure 7: Visualization of feature response to “Damaged” and “Destroyed” categories in different layers of deep models over three event cases. (a) Goma-VE-2021. (b) Beirut-EP-2023. (c) Libya-FL-2023. (d) Hawii-WF-2023. (e) Turkey-EQ-2023. (f) Acapulco-HC-2023. In the visualization, closer to red indicates larger response values, and closer to dark blue the opposite. See Appendix A for implementation details.
To better understand model behavior beyond performance metrics, we explore the internal attention patterns of trained models using class activation maps (CAMs). Figure 7 presents the CAM responses of ChangeOS across six representative disaster events, including volcano, explosion, flood, wildfire, earthquake, and hurricane. We observe that the attention distribution varies across disaster types. Taking Goma-VE-2021 as the example, for the “Destroyed” category, the encoder processing the optical imagery exhibits strong activations across nearly all built-up regions, accurately localizing individual buildings. This suggests the model has effectively learned to extract detailed structural cues from pre-event optical imagery. In contrast, the encoder handling the SAR imagery shows intense activation over the lava-covered region on the left, indicating that the model has identified this as a key signal for destruction—likely due to the significant backscatter changes caused by lava flow. In the “Damaged” category, however, the activations are more subtle. Attention is primarily focused near the boundary of the lava flow, where partial or ambiguous structural changes occur. In the decoder stage, the destroyed buildings in the lava-affected area are strongly activated, aligning well with the reference labels. Conversely, for the “Damaged” class, only a few regions are activated. This suggests that inferring partial damage from SAR imagery in volcanic disaster scenarios remains a significant challenge, as the subtle structural degradation is often not clearly reflected in SAR backscatter or texture.
Figure 8: IoU distribution of deep models over seven disaster types.
[…]
@@ 4.2 What have models learned and what can they learn?
In summary, these findings confirm both the promise and limitations of Optical–SAR combinations for all-weather, global-scale disaster response. While the combination performs well in events characterized by large-scale surface disruption (e.g., wildfires, volcanoes), it struggles with subtle or localized damage patterns. Incorporating richer data sources, such as fully polarimetric SAR and LiDAR, can further enhance the accuracy and reliability of future building damage assessments.
@@
Table 7: Further contributions to mIoU from post-processing algorithms. ChangeMamba (Chen et al., 2024) is used here as the baseline. Details on these algorithms are provided in Appendix D.
@@
[…]
4.3 Effect of post-processing methods
Post-processing techniques help refine the raw predictions from DL models, reducing noise, improving consistency, and ensuring spatial coherence in damage maps (Zheng et al., 2021). Here, we further explore the impact of post-processing algorithms. Table 7 presents the effect of three post-processing techniques applied to ChangeMamba. The post-processing methods evaluated include test-time augmentation, object-based major voting, and model ensembling. Further details on these methods are provided in Appendix D.
[…]
In summary, post-processing techniques play a crucial role in improving damage proxy maps. Future work can further explore adaptive post-processing strategies tailored to specific disaster types to enhance prediction reliability in multimodal EO data contexts.
Table 8: The mIoU on different events for different DL models in zero-shot and one-shot cross-event transfer setup. The highest values are highlighted in purple, and the second-highest results are highlighted in teal.
@@ 4.3 Effect of post-processing methods
4.4 Evaluation results on cross-event transfer setup
@@ 4.4.2 Unsupervised domain adaptation and semi-supervised learning methods
Table 9: Evaluation results of some unsupervised domain adaptation methods on zero-shot cross-event transfer setup. DeepLabV3+ is used as the baseline here. Detailed mIoU on each event is listed in Table 14.
@@
[…]
Table 10: Evaluation results of some semi-supervised learning methods on one-shot cross-event transfer setup. DeepLabV3+ is used as the baseline here. Detailed mIoU on each event is listed in Table 15.
@@
[…]
In the context of cross-event transfer, UDA and SSL methods naturally emerge as promising strategies to bridge the domain gap between source and target disaster events. In the zero-shot setting, UDA methods aim to improve model generalization by aligning the source and target domains without requiring any target labels. In the one-shot setting, SSL methods further leverage a small number of labeled samples along with abundant unlabeled data from the target event, making them especially appealing for real-world disaster scenarios where rapid annotation is infeasible. However, it remains an open question whether these UDA and SSL methods, originally developed for natural image domains, can effectively handle the challenges of multimodal EO data in complex disaster scenarios. To examine the applicability of these methods, we evaluate several representative UDA and SSL algorithms using DeepLabV3+ as the baseline model. The results are reported in Tables 9 and 10, respectively.
[…]

### v4

@@ 2 Methodology
2.1 Problem statement
The objective of building damage assessment is to interpret EO data of areas affected by a disaster by generating a building damage map that reflects the extent of damage to buildings. To achieve this, two common approaches are typically employed. One is to directly treat the building damage assessment task as a single semantic segmentation task (Adriano et al., 2021; Gupta and Shah, 2021). In this approach, the pre- and post-event images are taken as inputs of the model, and then the final damage map is directly predicted. This process can be formalized as $\mathbf{Y}^{dam}=\mathcal{M}^{seg}(\mathbf{X}^{\mathrm{T}_{1}},\mathbf{X}^{\mathrm{T}_{2}})$, where $\mathbf{X}^{\mathrm{T}_{1}}$ is the pre-event imagery, $\mathbf{X}^{\mathrm{T}_{2}}$ is the post-event imagery, $\mathcal{M}^{seg}\left(\cdot\right)$ is a semantic segmentation model, $\mathbf{Y}^{dam}$ is the obtained damage map. In the context of this paper, $\mathbf{X}^{\mathrm{T}_{1}}$ is VHR optical imagery and $\mathbf{X}^{\mathrm{T}_{2}}$ is VHR SAR imagery.
The second adopts the task decoupling approach (Gupta et al., 2019; Zheng et al., 2021), which breaks down building damage assessment into two subtasks: the building localization task, i.e., separating the building from the background, and the damage classification task, i.e., focusing on the classification between different levels of damage. This approach can be formulated as $\mathbf{Y}^{loc}=\mathcal{M}^{loc}(\mathbf{X}^{\mathrm{T}_{1}})$ and $\mathbf{Y}^{clf}=\mathcal{M}^{clf}(\mathbf{X}^{\mathrm{T}_{1}},\mathbf{X}^{\mathrm{T}_{2}})$, where $\mathbf{Y}^{loc}$ is the building localization map, $\mathbf{Y}^{clf}$ is the damage classification map, $\mathcal{M}^{loc}\left(\cdot\right)$ and $\mathcal{M}^{clf}\left(\cdot\right)$ are models for building localization and damage classification, respectively. $\mathcal{M}^{loc}\left(\cdot\right)$ and $\mathcal{M}^{clf}\left(\cdot\right)$ can be two separate models (Gupta et al., 2019) or a unified multi-task learning model (Zheng et al., 2021; Chen et al., 2022a, 2024). The final building damage map is obtained by combining the two outputs using a simple mask operation: $\mathbf{Y}^{dam}=\mathbf{Y}^{loc}\odot\mathbf{Y}^{clf}$. Since this work aims not only to provide a large-scale multimodal dataset to support all-weather disaster response, but also to offer insights for designing appropriate methods in future research, both approaches are employed in the experiments to compare their results.
It is worth noting that in this work, we focus on the formulation of building damage assessment as a bi-temporal task, where both pre- and post-event images are used as inputs. This formulation aligns closely with generic change detection tasks, which aim to identify changes between two time points. Conceptually, building damage assessment can be viewed as a specialized “one-to-many” semantic change detection problem (Zheng et al., 2021, 2024; Lu et al., 2024), where the objective is not only to detect whether a change has occurred but also to categorize the type and severity of changes (damages) to buildings. Many existing methods are thus derived from or adapted versions of generic change detection frameworks (Chen et al., 2024; Zheng et al., 2024; Guo et al., 2024).
@@ 2.2 Benchmark suites
Several advanced deep network architectures from both the computer vision and EO communities are evaluated on Bright. Since building damage assessment can be considered a specialized semantic segmentation task, we adopted two well-known segmentation networks from the computer vision field: UNet (Ronneberger et al., 2015) and DeepLabV3+ (Chen et al., 2018); and five state-of-the-art networks from the EO community: SiamAttnUNet (Adriano et al., 2021), SiamCRNN (Chen et al., 2020), ChangeOS (Zheng et al., 2021), DamageFormer (Chen et al., 2022a), and ChangeMamba (Chen et al., 2024). These seven networks encompass a broad range of representative DL architectures, including convolutional neural networks (CNNs), recurrent neural networks (RNNs), Transformers, and the more recent Mamba architecture. Among the seven networks, UNet, DeepLabV3+, and SiamAttnUNet adopt the first approach defined in Section 2.1, i.e., directly treating building damage assessment as a single semantic segmentation task. In contrast, SiamCRNN, ChangeOS, DamageFormer, and ChangeMamba adopt the second approach by decoupling the task into building localization and damage classification tasks.
Beyond supervised DL models, Bright also enables the evaluation of other learning strategies and methods commonly explored in the EO and computer vision communities:
@@
Table 5: Accuracy assessment for different DL models on the test set under the standard ML data split (set-level mIoU). The highest values are highlighted in purple, and the second-highest results are highlighted in teal.
Method $F_{1}^{loc}$ ($\%$) $F_{1}^{clf}$ ($\%$) Final OA ($\%$) Final mIoU ($\%$) IoU per class ($\%$)
@@ 2.3 Model training
To train the supervised models, we use a combination of cross-entropy loss and Lovasz softmax loss (Berman et al., 2018). Cross-entropy loss serves as the basic loss function for dense prediction tasks, while Lovasz softmax loss effectively addresses sample imbalance between non-building and building pixels and across different damage levels. For UNet, DeepLabV3+, and SiamAttnUNet, which directly predict damage maps from the input multimodal image pairs, the training loss function is defined as:
@@
2.4 Accuracy assessment
We adopt overall accuracy (OA), F1 score (F1), and mean intersection over union (mIoU) to evaluate the performance of the models. These are commonly used metrics in building damage assessment (Zheng et al., 2021). Following the setup in previous unimodal building damage assessment studies and the related xView2 Challenge (Gupta et al., 2019), the F1 score is used to assess the performance of the models in the building localization and damage classification subtasks. OA and mIoU are used to measure the overall quality of the building damage map, providing a comprehensive assessment of the models’ ability to localize buildings and classify damage levels accurately.
3 Results and Analysis
3.1 Evaluation on standard machine learning data split
[…]
@@ 3.1 Evaluation on standard machine learning data split
Figure 6: Damage maps predicted by different models on the test images of the 14 disaster events. The meaning of the color in reference maps and damage maps is consistent with Figure 4. The sources of EO images are illustrated in Table 2.
3.2 What have the models learned and what can they learn?
Figure 7: Visualization of feature response to “Damaged” and “Destroyed” categories in different layers of deep models over three event cases. (a) Goma-VE-2021. (b) Beirut-EP-2023. (c) Libya-FL-2023. (d) Hawii-WF-2023. (e) Turkey-EQ-2023. (f) Acapulco-HC-2023. In the visualization, closer to red indicates larger response values, and closer to dark blue the opposite. See Appendix B for implementation details.
To better understand the models’ behavior beyond performance metrics, we explore the internal attention patterns of the trained ChangeOS using class activation maps (CAMs) (Selvaraju et al., 2017). Figure 7 presents the CAM responses of ChangeOS across six representative disaster events: volcano, explosion, flood, wildfire, earthquake, and hurricane. We observe that the attention distribution varies across disaster types. Taking Goma-VE-2021 as an example, for the “Destroyed” category, the encoder exhibits strong activations in nearly all the built-up regions in the optical images, accurately localizing individual buildings. This suggests that the model has effectively learned to extract detailed structural cues from pre-event optical imagery. In contrast, for the SAR images, the encoder shows intense activation over the lava-covered regions on the left. This indicates that the model has identified the lava-covered regions as a key signal for destruction, likely due to the significant backscatter changes caused by lava flow. In the “Damaged” category, the activations are more subtle. Attention is primarily focused near the boundary of the lava flow, where partial or ambiguous structural changes occur. In the decoder, the destroyed buildings in the lava-affected area are strongly activated, which aligns well with the reference labels. Conversely, for the “Damaged” class, only a few regions are activated. This suggests that inferring partial damage from SAR imagery in volcanic disaster scenarios remains a significant challenge, as subtle structural degradation is often not clearly reflected in SAR backscatter or texture.
Figure 8: IoU distribution of deep models over seven disaster types. Each bar represents the average IoU of seven DL models for that specific category under each disaster type. The error bars indicate the standard deviation of IoU scores across the seven models.
Figure 9: Typical failure cases of different models on Bata-Explosion-2021 and Noto-Earthquake-2024 in Bright, where optical images are from © Maxar and © GSI Japan and SAR images are from © Capella Space and © Umbra.
@@ 3.2 What have the models learned and what can they learn?
These quantitative limitations are vividly illustrated by the typical failure cases shown in Figure 9. In the Bata-Explosion-2021 event, models misclassify severely destroyed buildings as intact, reflecting the difficulty of interpreting heterogeneous debris patterns. Similarly, in the Noto-Earthquake-2024 event, large-scale collapses are largely missed, highlighting the challenge of diverse and subtle seismic damage. These examples visually confirm that the significant heterogeneity in damage patterns makes it challenging for models to learn a consistent and generalizable representation of damage.
@@
In summary, these findings confirm both the promise and limitations of optical-SAR modality for all-weather, global-scale disaster response. Although this combination performs well in events characterized by large-scale surface disruption (e.g., wildfires, volcanoes), it struggles with subtle or localized damage patterns. Incorporating richer data sources, such as fully polarimetric SAR and LiDAR data, can further enhance the accuracy and reliability of future all-weather building damage assessments.
Table 7: Performance comparison of UNet and DeepLabV3+ using only post-event SAR input and pre-event optical plus post-event SAR inputs for damage classification task. Here, accurate building masks are provided as the post-processing step to all models to isolate the effect of building localization task on the damage classification task.
Method Modality $F_{1}^{clf}$ ($\%$) Final mIoU ($\%$) IoU per class ($\%$)
Background Intact Damaged Destroyed
UNet Post-event SAR 68.71 69.84 100.0 88.19 35.83 55.35
Pre-event optical + post-event SAR 73.59 72.41 100.0 89.38 44.83 55.42
DeepLabV3+ Post-event SAR 72.12 72.19 100.0 89.59 39.63 59.54
Pre-event optical + post-event SAR 73.90 73.93 100.0 90.32 40.45 64.94
Table 8: Performance comparison of different post-event modalities on a subset of Bright. Results are reported for UNet, DeepLabV3+, and DamageFormer on five disaster events where high-quality post-event optical imagery is available: Bata-Explosion-2020, Beirut-Explosion-2021, Hawaii-Wildfire-2023, Libya-Flood-2023, and Noto-Earthquake-2024.
Method Post-event modality $F_{1}^{loc}$ ($\%$) $F_{1}^{clf}$ ($\%$) Final mIoU ($\%$) IoU per class ($\%$)
Background Intact Damaged Destroyed
UNet SAR 85.05 71.43 62.94 94.39 65.60 42.34 49.43
Optical 86.46 75.64 65.96 94.56 68.62 45.27 55.36
Optical+SAR 86.70 74.46 66.29 94.72 69.16 41.73 59.57
DeepLabV3+ SAR 83.55 67.52 60.57 93.86 65.62 35.64 47.16
Optical 85.79 74.39 64.87 94.33 67.90 44.31 52.94
Optical+SAR 85.90 74.87 65.84 94.48 69.60 44.68 54.60
DamageFormer SAR 88.41 73.43 65.56 95.30 70.62 41.31 55.00
Optical 88.32 78.04 69.76 95.37 72.72 47.26 63.68
Optical+SAR 88.86 79.27 70.79 95.56 73.64 48.44 65.51
3.3 The role of optical pre-event data in multimodal building damage assessment
In the last section, CAM visualizations revealed that DL models also exhibit responses to disaster-specific patterns in pre-event optical imagery. This observation suggests that optical data may play a more complex role in multimodal building damage mapping than simply supporting building localization. In other words, in a multimodal bi-temporal setup, does pre-event optical imagery act solely as a localization aid, or does it provide additional semantic cues that networks can exploit for more accurate damage classification?
To explore this, we conducted controlled experiments using UNet and DeepLabV3+. Both networks were trained under two configurations: (i) using post-event SAR imagery only, and (ii) using multimodal pre- and post-event inputs (optical-SAR). To isolate the contribution of pre-event optical data beyond building localization, we provided perfect building masks for postprocessing in both settings. This design ensures that any observed differences in performance are attributable to the additional information from pre-event optical imagery, rather than differences in network architecture or localization accuracy.
The results, summarized in Table 7, show that incorporating pre-event optical imagery leads to notable improvements in distinguishing building damage levels. For UNet, the IoU for the “Damaged” class increased from 35.83% (SAR only) to 44.83% (Optical-SAR), and for the “Destroyed” class from 55.35% to 55.42%. DeepLabV3+ exhibited significant gains also, with IoU improvements from 39.63% to 40.45% for “Damaged” category, and from 59.54% to 64.94% for “Destroyed” category. These results suggest that pre-event optical imagery contributes beyond mere building localization, enriching the feature space for more effective semantic comparison for different building damage levels across modalities.
3.4 Impact of post-event modality on building damage assessment performance
Although the primary design of Bright is to facilitate all-weather disaster response through the use of pre-event optical and post-event SAR imagery, it is also important to understand how these modalities compare when high-quality post-event optical imagery is available. To this end, we conducted supplementary experiments on a subset of events, including Bata-Explosion-2020, Beirut-Explosion-2021, Hawaii-Wildfire-2023, Libya-Flood-2023, and Noto-Earthquake-2024, for which pre-processed post-event optical data were accessible. We evaluated three experimental setups: (i) optical-only (pre-event optical + post-event optical), (ii) SAR-only (pre-event optical + post-event SAR, i.e., the standard Bright setting), and (iii) optical+SAR fusion (pre-event optical + post-event optical + post-event SAR).
Table 8 presents the experimental results. As expected, when ideal post-event optical imagery is available, the optical-only setup achieves higher performance than the SAR-only setup. For example, with DamageFormer, the optical-only configuration reaches a final mIoU of 69.76%, compared to 65.56% for SAR-only. Importantly, the performance gap between optical and SAR is not substantial, demonstrating that SAR alone provides a strong alternative in the absence of usable optical imagery. Moreover, the fusion of optical and SAR consistently yields the best results across all tested models. For instance, DamageFormer’s mIoU further increases to 70.79% with Optical+SAR fusion, indicating that SAR contributes complementary information that strengthens performance even under optimal optical conditions.
These findings underscore two important insights. First, multimodal fusion is beneficial even when high-quality optical data are available, as SAR provides unique structural information that enriches the optical signal. Second, the performance of the SAR-only approach, being reasonably close to the optical-only results, highlights the practical value of SAR in real-world disaster scenarios where post-event optical imagery is often unavailable. Bright is therefore designed to advance the development of models for these realistic, often non-ideal, but operationally critical all-weather disaster response settings.
3.5 Effect of post-processing method
@@
Table 9: Further contributions to mIoU from post-processing algorithms. ChangeMamba (Chen et al., 2024) is used here as the baseline. Details on these algorithms are provided in Appendix E.
@@
Test-time augmentation 68.50 $\uparrow\!$ 0.87 51.95 $\uparrow\!$ 0.56
Object-based majority voting 67.22 $\downarrow\!$ 0.41 52.08 $\uparrow\!$ 0.69
Ensembling multiple models 68.45 $\uparrow\!$ 0.82 52.14 $\uparrow\!$ 0.75
All 68.86 $\uparrow\!$ 1.23 52.31 $\uparrow\!$ 0.92
@@
Table 10: The mIoU on different events for different DL models in zero-shot and one-shot cross-event transfer setups. The highest values are highlighted in purple, and the second-highest results are highlighted in teal.
@@
3.6 Evaluation on cross-event transfer setup
@@ 3.6.2 Unsupervised domain adaptation and semi-supervised learning methods
Table 11: Results of unsupervised domain adaptation methods adopted for zero-shot cross-event transfer setup. DeepLabV3+ (Chen et al., 2018) is used as the baseline here. Detailed mIoU on each event is listed in Table 16.
@@
AdaptSeg (Tsai et al., 2018) 36.05 $\uparrow\!$ 0.45
AdvEnt (Vu et al., 2019) 35.43 $\downarrow\!$ 0.17
CLAN (Luo et al., 2019) 36.01 $\uparrow\!$ 0.41
PyCDA (Lian et al., 2019) 33.19 $\downarrow\!$ 2.41
FDA (Yang and Soatto, 2020) 36.29 $\uparrow\!$ 0.69
Table 12: Results of semi-supervised learning methods adopted for one-shot cross-event transfer setup. DeepLabV3+ (Chen et al., 2018) is used as the baseline here. Detailed mIoU on each event is listed in Table 17.
@@
MT (Tarvainen and Valpola, 2017) 40.00 $\uparrow\!$ 1.18
CCT (Ouali et al., 2020) 39.49 $\uparrow\!$ 0.67
GCT (Ke et al., 2020) 39.34 $\uparrow\!$ 0.52
CPS (Chen et al., 2021) 37.06 $\downarrow\!$ 1.76
In the context of cross-event transfer, the UDA and SSL methods naturally emerge as promising strategies to bridge the domain gap between the source and the target disaster events. In the zero-shot setting, UDA methods aim to improve model generalization by aligning the source and target domains without requiring any target labels. In the one-shot setting, SSL methods leverage a small number of labeled samples with abundant unlabeled data from the target event, making them especially appealing for real-world disaster scenarios where rapid and comprehensive annotation is infeasible. It remains an open question whether these UDA and SSL methods, originally developed for natural image domains, can effectively handle the challenges of multimodal EO data in complex disaster scenarios. We evaluated several representative UDA and SSL methods using DeepLabV3+ as the baseline model to examine their capabilities in complex disaster scenarios. The results for UDA and SSL are reported in Tables 11 and 12, respectively.